40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Nov 18 01:51:37 2011, Koji, Update, IOO, Stochmon update stochmon.pdfstochmon_calib.pdfRFAM_PD_noise.pdf
    Reply  Fri Nov 25 20:03:05 2011, Koji, Update, IOO, Stochmon update stochmon_calib.pdf
       Reply  Mon Nov 28 18:19:53 2011, kiwamu, Update, IOO, Stochmon seems working RFAM_48hours.png
Message ID: 5941     Entry time: Fri Nov 18 01:51:37 2011     Reply to this: 6009
Author: Koji 
Type: Update 
Category: IOO 
Subject: Stochmon update 

Update of the stochmon status

[Attachment 1: Circuit diagram]

- The new stochmon has a low noise amplifier (MAR-6SM) inside.
The RFAM signal from the PD has the power of -60~-50dBm, which is almost at the bottom of the sensitivity for the power detector.

- The band pass filters were doubled.

- I've suffered from some RF coupling from the power line as the power detector is quite sensitive to it.
The situation has been largely improved by the EMI filters in the power supply path, although the problem is still present.
The worst remaining problem is that we can not close the aluminum lid as it cause a huge sprious coupling.

 

[Attachment 2: Calibration result]

- The outputs were calibarted with Marconi. They showed the signals linear to dBm for the input powers between -70dBm and -10dBm.

- The calibration result was fitted with the empirical fit function. The function and the results are shown in the attachment.

[Attachment 3: Detection limit]

- The attached figure shows the power spectrum of the PD output. This measurement gives us the amount of the RF power given from the PD noise when there is no RF signal.

11MHz out passband noise: −72.7dBm ===> V11 = 2.0483
30MHz out passband noise: −64.6dBm ===> V30 = 1.9333
55MHz out passband noise: −71.2dBm ===> V55 = 2.0272

- Now 11MHz and 55MHz outputs seem indicating the power correctly, but the 29.5MHz output never provides useful information.
It is a constant value independent from the state of the incident beam. Strangely this problem disappears if the marconi is used
for the RF source. Thus this issue is not seen in the calibration measurement.

- So far, 11MHz, 29.5MHz, 55MHz, and DC outputs appear in the channels C1:IOO-RFAMPD_33MHZ, C1:IOO-RFAMPD_133MHZ, C1:IOO-RFAMPD_166MHZ, and C1:IOO-RFAMPD_199MHZ.
They will be renamed.

Attachment 1: stochmon.pdf  55 kB  | Hide | Hide all
stochmon.pdf
Attachment 2: stochmon_calib.pdf  33 kB  | Hide | Hide all
stochmon_calib.pdf
Attachment 3: RFAM_PD_noise.pdf  28 kB  Uploaded Fri Nov 18 04:37:36 2011  | Hide | Hide all
RFAM_PD_noise.pdf
ELOG V3.1.3-